# Verifying the Resonance Schemes of Unstable Particles at Lepton Colliders

### v1

@@ Abstract
$\mathcal{D}_{\rm BW}\sim\frac{1}{p^{2}-m^{2}_{\rm BW}+im_{\rm BW}\Gamma_{\rm BW}},$ (1)
whose magnitude squared $1/[(p^{2}-m^{2}_{\rm BW})^{2}+m^{2}_{\rm BW}\Gamma^{2}_{\rm BW}]$ follows a resonance shape with pole at exactly the mass, $p^{2}=m^{2}_{\rm BW}$. The Breit-Wigner (BW) scheme give a very intuitive understanding of the pole mass $m_{\rm BW}$. Since the propagator can enter concrete processes in various ways, we refrain from replacing $p^{2}$ by the Mandelstam variable $s$ that is usually for the $s$-channel processes to make it general. A review of more details can be found in Giacosa:2021mbz .
A physical particle should have definite dispersion relation, $E=\sqrt{\bm{p}^{2}+m^{2}}$ where $(E,\bm{p})$ is the 4-momentum and $m$ is mass, according to Einstein’s special relativity. With a real energy $E$, the particle state evolves into the infinite future with a time-dependent evolution phase $e^{-iEt}$. So a real energy $E$ associated with time evolution intrinsically describes a stable particle. On the other hand, an unstable particle develops an imaginary part $-\frac{i}{2}\Gamma$ that is attached to the real energy, $E\rightarrow E-\frac{i}{2}\Gamma$. Then the particle wave function $\sim e^{-iEt-\frac{1}{2}\Gamma t}$ keeps decreasing with time and the corresponding decay width (lifetime) is $\Gamma$ ($1/\Gamma$).
It is much more intuitive to define the exact value of decay width in the rest frame. In other words, the physical pole should actually happen at $E=m-\frac{i}{2}\Gamma\equiv\mu$. Considering the fact that a relativistic dispersion relation should be Lorentz invariant, the propagator contains $p^{2}$ in the denominator Willenbrock:1991hu ; Willenbrock:2022smq ,
@@
The pole $E=\sqrt{\bm{p}^{2}+\mu^{2}}$ then happens at $\mu$ in the rest frame with vanishing $\bm{p}$. Since this scheme comes from theoretical arguments and has the benefit of being gauge independent Sirlin:1991fd , we call it the theoretical (th) scheme.
Comparing with the Breit-Wigner resonance in Eq. (1), the theoretical scheme has an extra term $\frac{1}{4}\Gamma^{2}$. For a narrow resonance, $\Gamma\ll m$, the Breit-Wigner scheme is a good enough approximation. However, the $Z$ mass can shift by 8.5 MeV which seems to be already larger than the current precision $\Delta m_{Z}=2.1$ MeV Workman:2022ynf .
@@
The Breit-Wigner and energy-dependent schemes agree quite well with each other for the narrow width approximation ($\Gamma\ll m$). Taking the weak gauge bosons for illustration, the width is just 2.7% for both $Z$ ($m_{Z}=91.1876$ GeV vs $\Gamma_{Z}=2.4955$ GeV Workman:2022ynf ) and $W$ ($m_{W}=80.376$ GeV vs $\Gamma_{W}=2.195$ GeV Workman:2022ynf ). Consequently, the difference $\Gamma^{2}/2m^{2}$ between the two schemes is only at the level of $3.7\times 10^{-4}$ for $Z$ and $W$ corresponding to a shift of 34 MeV and 30 MeV in the $Z$ and $W$ masses, respectively. Although such correction seems quite tiny, the sensitivity at lepton colliders can reach the precision. For the $Z$ boson mass, the current precision that mainly comes from LEP is 2.1 MeV while it can further shrink to 0.1 MeV at CEPC CEPCPhysicsStudyGroup:2022uwl and FCC-ee Blondel:2021ema . It seems possible and is of great importance to use the in-situ $Z$ lineshape scan to test the resonant schemes for an unambiguous conclusion.
@@
The only way for distinguishing the Breit-Wigner and theoretical schemes probably lies in making theoretical connection between $m_{Z}$ and $\Gamma_{Z}$. As shown in Eq. (10), the mass and decay width are inversely proportional to each other. When extracting the decay width from data, the decay width takes a smaller value for a heavier $Z$ boson, $\Gamma_{Z}\propto 1/m_{Z}$, using either the Breit-Wigner or the theoretical scheme. However, the naive expectation with dimensional analysis tells us that the decay width of an unstable particle is typically proportional to its mass, $\Gamma_{Z}\propto m_{Z}$ for the $Z\rightarrow f\bar{f}$ decay,
@@
This mismatch between the extracted $\Gamma_{Z}$ from data and its theoretically predicted value can be experimentally tested. To be more specific, once the $Z$ lineshape fit is done with the Breit-Wigner scheme for example, the best-fit value of $m_{\rm BW}$ can be used to predict $\Gamma_{Z}$ as Eq. (11). Since there is no real data yet, the central values are taken as the current $m_{Z}=91.1876$ GeV Workman:2022ynf and the derived $\Gamma_{Z}=2.47042$ GeV with Eq. (11). For such setup, the Breit-Wigner scheme with one fitting parameter $m_{\rm BW}$ can perfectly fit the pseudo-data. For the theoretical scheme, we take $m_{\rm th}$ as the fitting parameter and $\Gamma_{\rm th}(m_{\rm th})$ is consequently predicted as function of $m_{\rm th}$ according to Eq. (11). However the fitting should still be be done after conversion to the Breit-Wigner scheme,
@@
Comparing with the $9.2\times 10^{-5}$ correction of Eq. (10), a thumb rule is that both precisions $\Delta m_{Z}\lesssim 8.4$ MeV and $\Delta\Gamma_{Z}\lesssim 0.023$ MeV are necessary to have sizable sensitivity. The current sensitivities $\Delta m_{Z}=2.1$ MeV and $\Delta\Gamma_{Z}=2.3$ MeV Workman:2022ynf are not good enough. However, with the projected sensitivity $\Delta m_{Z}=0.1$ MeV and $\Delta\Gamma_{Z}=0.025$ MeV at CEPC CEPCPhysicsStudyGroup:2022uwl and FCC-ee Blondel:2021ema , $\chi^{2}_{Z,\rm min}$ can reach 326. For the ILC-GigaZ projection of $\Delta m_{Z}=0.2$ MeV and $\Delta\Gamma_{Z}=0.12$ MeV Belloni:2022due , the projected sensitivity can also touch $\chi^{2}_{Z,\rm min}=14.3$.
@@
$+\left[\frac{G_{F,\rm th}-G_{F}}{\Delta G_{F}}\right]^{2},$ (16)
where $R_{W}$ takes the similar form as $R_{Z}$ in Eq. (10). For the central values we take the current $m_{W}=80.376$ GeV Workman:2022ynf and the derived $\Gamma_{W}=2.04501$ from Eq. (15), in addition to $G_{F}=1.166\,378\,8\times 10^{-5}\,\mbox{GeV}^{-2}$ Workman:2022ynf . The weak coupling is then fixed to $g=0.652986$ by Eq. (13) in the Breit-Wigner scheme. The sensitivity can reach $\Delta\chi^{2}_{W,\rm min}=169$, 263, and 7.3 at CEPC, FCC-ee, and ILC, respectively. By fitting the $WW$ threshold scan data, the extract $W$ mass values from the Breit-Wigner and theoretical schemes should also follow the scheme conversion in Eq. (10). However, the precisely measurement Fermi constant tends to give exactly the same value for $m_{W}$ in both schemes. This contradiction gives the distinguishing power.

### v2

@@ Abstract
$\mathcal{D}_{\rm BW}\sim\frac{1}{p^{2}-m^{2}_{\rm BW}+im_{\rm BW}\Gamma_{\rm BW}}$ (1)
in natural units ($\hbar=c=1$), whose magnitude squared $1/[(p^{2}-m^{2}_{\rm BW})^{2}+m^{2}_{\rm BW}\Gamma^{2}_{\rm BW}]$ follows a resonance shape with pole at exactly the mass, $p^{2}=m^{2}_{\rm BW}$. The Breit-Wigner (BW) scheme give a very intuitive understanding of the pole mass $m_{\rm BW}$. Since the propagator can enter concrete processes in various ways, we refrain from replacing $p^{2}$ by the Mandelstam variable $s$ that is usually for the $s$-channel processes to make it general. A review of more details can be found in Bohm:2004zi ; Giacosa:2021mbz .
A physical particle should have definite dispersion relation, $E=\sqrt{\bm{p}^{2}+m^{2}}$ where $(E,\bm{p})$ is the 4-momentum and $m$ is mass. With a real energy $E$, the particle state evolves into the infinite future with a time-dependent evolution phase $e^{-iEt}$. So a real energy $E$ associated with time evolution intrinsically describes a stable particle. On the other hand, an unstable particle develops an imaginary part $-\frac{i}{2}\Gamma$ that is attached to the real energy, $E\rightarrow E-\frac{i}{2}\Gamma$. Then the particle wave function $\sim e^{-iEt-\frac{1}{2}\Gamma t}$ keeps decreasing with time and the corresponding decay width (lifetime) is $\Gamma$ ($1/\Gamma$).
It is much more intuitive to define the exact value of decay width in the rest frame. In other words, the physical pole should actually happen at $E=m-\frac{i}{2}\Gamma\equiv\mu$, which is consistent with the linear approximation for the non-relativistic limit Dalitz:1961dx ; Dalitz:1963xk . Considering the fact that a relativistic dispersion relation should be Lorentz invariant, the propagator contains $p^{2}$ in the denominator Willenbrock:1991hu ; Willenbrock:2022smq ,
@@
The pole $E=\sqrt{\bm{p}^{2}+\mu^{2}}$ then happens at $\mu$ in the rest frame with vanishing $\bm{p}$. Since this scheme comes from theoretical arguments and has the benefit of being gauge independent Sirlin:1991fd ; Papavassiliou:1995fq , we call it the theoretical (th) scheme.
Comparing with the Breit-Wigner resonance in Eq. (1), the theoretical scheme has an extra term $\frac{1}{4}\Gamma^{2}$. For a narrow resonance, $\Gamma\ll m$, the Breit-Wigner scheme is a good enough approximation. By solving the equations, $m^{2}_{\rm BW}=m^{2}_{\rm th}-\frac{1}{4}\Gamma^{2}_{\rm th}$ and $m_{\rm BW}\Gamma_{\rm BW}=m_{\rm th}\Gamma_{\rm th}$, however, the $Z$ mass can shift by $m_{\rm th}-m_{\rm BW}=8.5$ MeV which seems to be already larger than the current precision $\Delta m_{Z}=2.1$ MeV Workman:2022ynf .
@@
The Breit-Wigner and energy-dependent schemes agree quite well with each other for the narrow width approximation ($\Gamma\ll m$). Taking the weak gauge bosons for illustration, the width is just 2.7% of the mass for both $Z$ ($m_{Z}=91.1876\pm 0.0021$ GeV vs $\Gamma_{Z}=2.4955\pm 0.0023$ GeV at LEP 1 Workman:2022ynf ; ALEPH:2005ab ) and $W$ ($m_{W}=80.376\pm 0.033$ GeV vs $\Gamma_{W}=2.195\pm 0.83$ GeV at LEP 2 Workman:2022ynf ; ALEPH:2013dgf ). Consequently, the relative difference $\Gamma^{2}/2m^{2}$ between the two schemes is only at the level of $3.7\times 10^{-4}$ for the $Z$ and $W$ bosons. This corresponds to shifts of 34 MeV and 30 MeV in the $Z$ and $W$ masses, respectively, and correspondingly a shift of 0.93 MeV and 0.82 MeV in their decay widths. Although such correction seems quite tiny, the sensitivity at lepton colliders can reach the precision.  [^1] The recent measurements of $m_{W}$ and $\Gamma_{W}$ at the LHC achieved a precision of $\Delta m_{W}=15.9$ MeV and $\Delta\Gamma_{W}=47$ MeV from the ATLAS Collaboration ATLAS:2024erm . Similarly, the CMS Collaboration reached $\Delta m_{W}=9.9$ MeV CMS:2024nau while $\Delta\Gamma_{W}$ was not explicitly reported. To distinguish between the two resonance schemes, high precision in both the mass and decay width measurements is essential. Despite having achieved sufficient precision in mass measurements, it is still challenging for hadron colliders to differentiate between the resonance schemes. For the $Z$ boson mass, the current precision that mainly comes from LEP is 2.1 MeV while it can further shrink to 0.1 MeV at CEPC CEPCPhysicsStudyGroup:2022uwl and FCC-ee Blondel:2021ema . It seems possible and is of great importance to use the in-situ $Z$ lineshape scan to test the resonant schemes for an unambiguous conclusion.
@@
Notice that the soft real and radiative virtual corrections can have significant effect on the forward-backward asymmetry $A^{f}_{\rm FB}$ Greco:1980mh . Especially, the value of $A^{f}_{\rm FB}$ reduces above the $Z$ pole which would make the relative difference between different resonance schemes even larger. As a conservative estimation of the sensitivity, we adopt the tree-level calculation and leave the more sophisticated soft real and radiative virtual corrections for future study.
@@
The only way for distinguishing the Breit-Wigner and theoretical schemes probably lies in making extra connection between $m_{Z}$ and $\Gamma_{Z}$. Eq. (10) requires the decay width to be inversely proportional to the mass, $\Gamma_{Z}\propto 1/m_{Z}$. This relationship arises because the observable quantity as data is the imaginary part of the self-energy function, ${\rm Im}\,\Pi_{Z}=m_{Z}\Gamma_{Z}$. When extracting the decay width from data, a heavier $Z$ boson corresponds to a smaller decay width, $\Gamma_{Z}\propto 1/m_{Z}$, regardless whether the Breit-Wigner or the theoretical scheme is used. However, the naive expectation with dimensional analysis tells us that the decay width of an unstable particle is typically proportional to its mass, $\Gamma_{Z}\propto m_{Z}$ for the $Z\rightarrow f\bar{f}$ decay,
@@
This mismatch between the extracted $\Gamma_{Z}$ from data and its theoretically predicted value can be experimentally tested. To be more specific, once the $Z$ lineshape fit is done with the Breit-Wigner scheme for example, the best-fit value of $m_{\rm BW}$ can be used to predict $\Gamma_{Z}$ as Eq. (11). Since there is no real data yet, the central values are taken as the current $m_{Z}=91.1876$ GeV Workman:2022ynf and the derived $\Gamma_{Z}=2.47042$ GeV with Eq. (11). For such setup, the Breit-Wigner scheme with one fitting parameter $m_{\rm BW}$ can perfectly fit the pseudo-data. For the theoretical scheme, we take $m_{\rm th}$ as the fitting parameter and $\Gamma_{\rm th}(m_{\rm th})$ is consequently predicted as function of $m_{\rm th}$ according to Eq. (11). However the fitting should still be done after conversion to the Breit-Wigner scheme,
@@
Comparing with the $9.2\times 10^{-5}$ correction of Eq. (10), a thumb rule is that both precisions $\Delta m_{Z}\lesssim 8.4$ MeV and $\Delta\Gamma_{Z}\lesssim 0.23$ MeV are necessary to have sizable sensitivity. The current sensitivities $\Delta m_{Z}=2.1$ MeV and $\Delta\Gamma_{Z}=2.3$ MeV Workman:2022ynf are not good enough. However, with the projected sensitivity $\Delta m_{Z}=0.1$ MeV and $\Delta\Gamma_{Z}=0.025$ MeV at CEPC CEPCPhysicsStudyGroup:2022uwl and FCC-ee Blondel:2021ema , $\chi^{2}_{Z,\rm min}$ can reach 326. For the ILC-GigaZ projection of $\Delta m_{Z}=0.2$ MeV and $\Delta\Gamma_{Z}=0.12$ MeV Belloni:2022due , the projected sensitivity can also touch $\chi^{2}_{Z,\rm min}=14.3$. We have also tried implementing the energy-dependent width for both the Breit-Wigner and theoretical schemes. Then the $Z$ lineshape scan should be used to give $\Delta\chi^{2}_{\rm min}=349$ and $696$ at CEPC and FCC-ee, respectively.
@@
$+\left(\frac{G_{F,\rm th}-G_{F}}{\Delta G_{F}}\right)^{2},$ (16)
where $R_{W}$ takes the similar form as $R_{Z}$ in Eq. (10). For the central values we take the current $m_{W}=80.376$ GeV Workman:2022ynf and the derived $\Gamma_{W}=2.04501$ from Eq. (15), in addition to $G_{F}=1.166\,378\,8\times 10^{-5}\,\mbox{GeV}^{-2}$ Workman:2022ynf . The weak coupling is then fixed to $g=0.652986$ by Eq. (13) in the Breit-Wigner scheme. The sensitivity can reach $\Delta\chi^{2}_{W,\rm min}=169$, 263, and 7.3 at CEPC, FCC-ee, and ILC, respectively. By fitting the $WW$ threshold scan data, the extracted $W$ mass values from the Breit-Wigner and theoretical schemes should also follow the scheme conversion in Eq. (10). However, the precise measurement of the Fermi constant tends to give exactly the same value for $m_{W}$ in both schemes. These different scaling behaviors provide the distinguishing power.